OpenStudy (anonymous):

Prove that the convergence of ∑an (a sub-n) an implies the convergence of ∑√(an)/n if an > 0

OpenStudy (anonymous):

Let\[\sum a_n\]be a convergent sequence. Then it is bounded. Moreover, by the Bolzano-Weierstrass theorem, it follows that since\[\sum_{n=1}^{\infty}\frac{\sqrt{a_n}}{n}\]is a subsequence of the above one, it converges as well.